# Bridging the Synthesizability Gap in Perovskites by Combining Computations, Literature Data, and PU Learning

### v1

@@ 1 Introduction
Perovskites and their derivatives have become increasingly prominent for a variety of applications over the last couple of decades. Due to their tunable band gaps (0.5 eV to 6 eV) [1, 2], high light absorption coefficients ($10^{5}$ cm-1) [3], and low cost of synthesis, they are highly sought-after commercial replacements for Si as solar absorbers. ABX3 halide perovskites (HaPs) have been the most studied subclass of materials for photovoltaics [4] and photocatalysis [5]. Among them, hybrid organic-inorganic perovskites, which have an organic cation such as methylammonium or formamidinium at the A-site, have shown the best photovoltaic efficiencies [6], but have also suffered from thermal and mechanical instability [7, 8]. Hybrid perovskites can undergo protonation in the presence of moisture and exhibit undesired phase changes at elevated temperatures [9]. Inorganic perovskites do not have the same issues but typically contain Pb at the B-site and thus suffer from Pb toxicity [10].
@@
The many challenges of halide perovskites are being actively tackled via composition engineering [11, 12, 13, 14], including complex alloying at anion and/or cation sites [15, 16]. Much effort has recently been put into going beyond halide perovskites to chalcogenide perovskites (ChPs) [17, 18] and various perovskite derivatives [19, 20]. Compounds in many perovskite-derived chemical classes, such as A2BB’X6 double perovskites (DP), A2BX6 vacancy-ordered double perovskites (VO), and A3B2X9 compounds, effectively overcome some of the aforementioned issues. Due to the presence of such different types of perovskites, the exploration space is massive and complex. High-throughput density functional theory (HT-DFT) computations have been extensively used to perform screening across these chemical spaces [21, 22, 23] since it is infeasible to synthesize and characterize all possible compounds without prior prediction of their properties. DFT datasets help researchers comprehensively map the composition-structure-property space of perovskites to aid experimental studies.
@@
We have performed various perovskite screening studies in our research group over the past few years [24, 25, 26], leading to a large DFT dataset of ABX3 halide perovskites in a chemical space spanning 14 different A/B/X species and four prototype perovskite phases. This dataset includes inorganic and hybrid perovskites and structures with alloying at any of the cation or anion sites. It includes >> 1000 unique perovskite compositions with properties such as band gap, decomposition energy, and spectroscopic limited maximum Efficiency (SLME) [27] computed from various semi-local GGA and non-local HSE06 functionals [28, 29, 30, 31, 32, 33]. The functionals we used include GGA-PBE with van der Waals D3 corrections [30, 31] for hybrid perovskites, GGA-PBEsol for inorganic compounds, and both static and full relaxation HSE06 with and without spin orbit coupling (SOC, important for systems containing heavier atoms). Our recent work [25] highlights how different functionals affect the properties of interest for different types of perovskites, providing some understanding about the levels of theory necessary for reproducing experimental results. While a majority of our previous dataset is focused on halide perovskites, it has now been extended to chalcogenide perovskites and a variety of double perovskites.
@@
Given that our dataset is well-representative of the perovskite chemical space, it is suitable for training surrogate machine learning (ML) models [34, 35] which can then be deployed for accelerated prediction and screening across millions of possible compositions. Using the DFT data, we trained regularized greedy forest models to predict the band gap and decomposition energy (which shows whether a compound would readily decompose to alternative binary phases) of any ABX3 halide perovskite. This is different from energy above hull that takes into account all possible combinations of phases (theoretical and experimental). We choose decomposition energy as a possible stability metric because the primary decomposition pathway for inorganic and hybrid perovskites is to their binary phases [36], and most perovskite films are synthesized using these known binary precursors [37]. These properties are central in determining the stability and photovoltaic suitability of the materials. Since we have data from different levels of theory, we trained these models within a multi-fidelity framework [38] such that the source of data (or DFT functional) is part of the input. We used both compositional information and well-known elemental or molecular properties as part of the descriptors and achieved prediction accuracy of $\sim$ 98% for the properties of interest.
@@
The best regression models were used to make on-demand predictions for thousands of enumerated ABX3 compounds across the four prototype phases, leading to the discovery of (a) novel stable high-efficiency single-junction solar absorber candidates [38], and (b) novel candidates with high predicted solar-to-hydrogen (STH) efficiencies for photocatalytic water splitting [39]. The latter work utilized empirically predicted electronic band edges to place them relative to H2O redox potential levels. Importantly, the best materials found via DFT+ML screening were also divided into different categories, such as hybrid vs purely inorganic, and Pb-free vs Pb-containing compounds. We recently extended this work to structure-based predictions using crystal graph neural network models [40, 41, 42]. By accumulating a massive perovskite dataset of $\sim$ 30,000 structures and their DFT-computed energies, atomic forces, and stresses from across all our recent publications, we “fine-tuned” the M3GNET model [43] to develop a comprehensive interatomic potential that is an effective DFT surrogate for our chosen perovskite chemical space [44]. Our rigorously optimized “Perovs-IAP” model makes accurate energy and force predictions for bulk, surface slab, and defect-containing perovskite structures, and is capable of rapid geometry optimization of any such structure.
@@
Thus, we now have a rich perovskite dataset and versatile descriptor-based and structure-based predictive models. However, DFT predictions of the decomposition energy or energy above hull are not entirely indicative of the likelihood of synthesis of any novel compound. In the past, the Goldschmidt tolerance factor [45] and octahedral factor [46] have been used to determine the general suitability of constituent ions to form a structurally robust perovskite based on their ionic radii, but satisfying these factors does not necessarily translate to a compound being synthesizable, especially when it comes to complex perovskite alloys with a distribution of anions and cations. DFT energies are often used for evaluating formability, but are clearly insufficient because calculations are performed at 0 K and do not account for entropic stabilization and disorder unless specifically included. What comes to the rescue here is the existence of many perovskite compounds that were actually synthesized and reported in the literature, and may inform the synthesis likelihood of other similar materials identified via computations.
@@
Classification techniques could be used for synthesis predictions provided the availability of sufficient positive and negative labels, but since studies only report positive examples, typical classifiers are not applicable. As material properties are emergent, if we can map their chemical descriptors to a synthesis label, we could, in theory, predict the order of synthesizability of different compositions. However, this requires a semi-supervised learning approach since it may be nearly impossible to classify a material as definitely non-synthesizable. Publications seldom report compositions that were attempted but failed to materialize in the lab. This means one could only apply positive labels to some of the materials and unknown labels to the remaining materials, which have not been synthesized to date but could be in the future. One technique that has recently been used by researchers for synthesis predictions is “Positive-Unlabeled (PU) learning” [47, 48, 49], where the classifier learns from the positive examples and assigns a probability to the unlabeled points based on similarity of the descriptors.
@@
Frey et al. [50] demonstrated the use of PU learning in predicting synthesis scores for MXenes, a novel class of layered materials. They used a decision tree classifier with specific MAX (the bulk phase from which MXenes are derived) and MXene features such as the number of layers, in-plane lattice constant, bond lengths, etc., which were combined with DFT-computed features such as formation energies and Bader charges. The labels for the synthesized compounds were obtained from their work on synthesizing MXenes and MAX phases. The model was able to predict 18 new MXene compositions as synthesizable out of 118 unlabeled samples. Another study by Jang et al. [51] used the PU learning approach but combined it with atomic features and the crystal-graph convolutional neural network (CGCNN) framework [52]. This work used the entire Materials Project (MP) [53] dataset and provided positive synthesis labels to structures having an Inorganic Crystal Structure Database (ICSD) [54] tag. Synthesis probabilities were predicted for materials in the Open Quantum Materials Database (OQMD) [55] and the MP-Virtual database; $\sim$ 12% of the total compounds from OQMD and $\sim$ 20% from MP-virtual datasets were found to be synthesizable.
@@
Another study on a similar dataset was performed by Gu et al. [56] who trained a graph convolution network based on MEGNet [57] using a combination of perovskite data from AFLOW [58], Materials Project, and OQMD [55]. Domain-specific transfer learning was performed to make the model accurate for perovskite structures. A recent study by Zhao et al. [59] utilized a similar approach for Metal Organic Frameworks (MOFs) and used a CGCNN-based classifier model. Compounds were divided based on pre-defined rules, following which their synthesizability scores were predicted. $\sim$ 12% of the compounds from across the unlabeled set were found to be synthesizable. Yet another recent study from Chung et al. [60] predicted the synthesizability of ternary oxides using PU learning, with labels applied specifically for oxides that can be synthesized via solid-state methods.
@@
A major consideration is, of course, the collection of synthesis labels for as many compounds as possible. To do this, we used large language models (LLMs) [61] and extracted chemical formulas of ABX3, A2BX6, or A2BB’X6 halide or chalcogenide compounds reported to be synthesized in various publications. We additionally applied labels to compounds based on their existence in the ICSD, ultimately resulting in 239 compounds with positive synthesis labels. This was used along with a dataset of 832 unlabeled compounds simulated using the GGA-PBE functional to train combined models for predicting both DFT properties and the synthesis probability. Based on prediction and screening across thousands of hypothetical compounds, we discovered potentially hundreds of new stable and synthesizable perovskites with promising optoelectronic properties. Our general workflow is presented in Figure 1. In the following sections, we describe the computational data generation, experimental label extraction, descriptors, different ML models, and results of training and prediction over unlabeled samples. We discuss the best candidates thus identified and present an outlook for the future of perovskite discovery.
@@ 2.1.1 Mannodi Group Dataset
DFT computations performed in our group over the years and available as part of multiple publications [24, 25, 26] have resulted in an extensive database of halide and chalcogenide (single and double) perovskite structures across cubic, tetragonal, orthorhombic, and hexagonal phases. A distribution of the ions that adopt the A, B/B’, and X sites in these compounds is pictured in Figure 2(a), with different colors referring to ABX3 halide perovskites (HaP), ABX3 chalcogenide perovskites (ChP), A2BB’X6 double perovskites (DP), and A2BX6 vacancy-ordered double perovskites (VO). As explained later, the identity and fractions of these A/B/B’/X species, along with their phases, are used to create perovskite feature vectors that are inputs to various ML models. A combinatorially massive number of A-B-X compositions could be created within the highlighted chemical space with arbitrary mixing at any site, and only a small fraction of it was ultimately used for creating the DFT dataset.
@@
Figure 2 (b) shows the computed decomposition energy (in eV per ABX3 formula unit) plotted against the band gap (in eV) for a total of 851 compounds, from GGA-PBE calculations. While our prior work also reports data from other GGA and HSE06 functionals, we focus on the PBE dataset for this work because of its relatively larger coverage of the perovskite chemical space of interest. Even though the majority of the dataset is cubic and HaP compositions, the models we report later perform well for all phases and ChP, VO, and DP compounds. The band gap values range from close to 0 eV to greater than 6 eV, and decomposition energies range from -1.5 eV (indicating high stability) to 1.5 eV. The formula to calculate the decomposition energy is shown in Eq. 1, where $E_{comp}$ stands for the total energy from DFT for the given composition, and the final term accounts for the mixing entropy in alloys. As an example, the decomposition energy for a compound with B-site mixing will be calculated as shown in Eq. 2. These formulas hold for both ABX3 HaP and ChP compounds. For A2BB’X6 double perovskites, we use binaries based on both B and B’ cations. For A2BX6 vacancy ordered double perovskites, the B site has an oxidation state of 4+, and thus we use BX4 as the binary phase. Most of the larger band gaps arise from double perovskites due to the lack of overlap between BX6 octahedra owing to the alternation of B and B’ cations in the structure. Decomposition energies tend to cluster near the 0 eV mark, and we usually consider a range of -0.5 eV to 0.5 eV as stable, keeping in mind theoretical uncertainties and possible entropic contributions. Other computational details can be found in past works [24, 25, 26].
@@ 2.1.2 Materials Project Data
We also queried the Materials Project (MP) [53] database to find perovskite structures distinct from the compounds in our dataset and subsequently combined all the compounds into one dataset. We first filtered all possible compounds with general formulas of ABX3, A2BB’X6, and A2BX6, where the anion X is either a halogen or chalcogen. The MP filtering workflow is pictured in Figure 3. This process led to a total of about 900 compounds. Then, we used the Robocrystallographer tool [62] to filter out the compounds that did not contain specific keywords such as “perovskite”, “perovskite-derived”, and “octahedral”, following which a manual examination was performed to verify that the parsed compound is indeed a perovskite structure. This finally led to 225 compounds, out of which 111 were already present in the Mannodi group dataset. For the 114 unique compounds, we performed additional DFT computations using the exact same GGA-PBE input parameters and convergence criteria used in our dataset, thus obtaining their band gaps and decomposition energies. Subsequently, compounds with zero band gaps were excluded from the training set to ensure model stability, resulting in 58 new compounds which were added to our original dataset to yield an expanded dataset of 909 points. We also used the “ICSD tag” (indicating experimental existence) available in the MP for any of the compounds to add synthesis labels to them. Out of the 225 compounds collected from MP, 33 had a valid ICSD tag, showing that they had been previously successfully synthesized. The computed properties pictured in Figure 2(b) are for the combined PBE dataset of 909 unique compounds.
Figure 3: The workflow used to extract valid perosvkite compounds from the MP [53] database; this data is then combined with our DFT dataset following quick calculations using uniform computational parameters.

### v2

@@ 1 Introduction
Perovskites and their derivatives have become increasingly prominent for a variety of applications over the last couple of decades. Due to their tunable band gaps (0.5 eV to 6 eV) [bandgap-1, bandgap-2], high light absorption coefficients ($10^{5}$ cm-1) [absorption-coeff], and low cost of synthesis, they are highly sought-after commercial replacements for Si as solar absorbers. ABX3 halide perovskites (HaPs) have been the most studied subclass of materials for photovoltaics [halide-photovol] and photocatalysis [halide-photocat]. Among them, hybrid organic-inorganic perovskites, which have an organic cation such as methylammonium or formamidinium at the A-site, have shown the best photovoltaic efficiencies [hybrid-1], but have also suffered from thermal and mechanical instability [hybrid-instability-1, hybrid-instability-2]. Hybrid perovskites can undergo protonation in the presence of moisture and exhibit undesired phase changes at elevated temperatures [hybrid-instability-thermal]. Inorganic perovskites do not have the same issues but typically contain Pb at the B-site and thus suffer from Pb toxicity [lead-toxicity-1].
@@
The many challenges of halide perovskites are being actively tackled via composition engineering [lead-free-1, composition-tuning-1, composition-tuning-2, stability], including complex alloying at anion and/or cation sites [alloying-1, alloying-2]. Much effort has recently been put into going beyond halide perovskites to chalcogenide perovskites (ChPs) [chalco-perovs, chalco-perovs-2] and various perovskite derivatives [derivative-1, derivative-2]. Compounds in many perovskite-derived chemical classes, such as A2BB’X6 double perovskites (DP), A2BX6 vacancy-ordered double perovskites (VO), and A3B2X9 compounds, effectively overcome some of the aforementioned issues. Due to the presence of such different types of perovskites, the exploration space is massive and complex. High-throughput density functional theory (HT-DFT) computations have been extensively used to perform screening across these chemical spaces [ht-dft-abse3, ht-dft-vo, ht-dft-dp] since it is infeasible to synthesize and characterize all possible compounds without prior prediction of their properties. DFT datasets help researchers comprehensively map the composition-structure-property space of perovskites to aid experimental studies.
@@
We have performed various perovskite screening studies in our research group over the past few years [htdft-1, htdft-2, htdft-3], leading to a large DFT dataset of ABX3 halide perovskites in a chemical space spanning 14 different A/B/X species and four prototype perovskite phases. This dataset includes inorganic and hybrid perovskites and structures with alloying at any of the cation or anion sites. It includes >> 1000 unique perovskite compositions with properties such as band gap, decomposition energy, and spectroscopic limited maximum Efficiency (SLME) [slme] computed from various semi-local GGA and non-local HSE06 functionals [pbe, pbesol, d3-1, d3-2, hse-1, hse-2]. The functionals we used include GGA-PBE with van der Waals D3 corrections [d3-1, d3-2] for hybrid perovskites, GGA-PBEsol for inorganic compounds, and both static and full relaxation HSE06 with and without spin orbit coupling (SOC, important for systems containing heavier atoms). Our recent work [htdft-2] highlights how different functionals affect the properties of interest for different types of perovskites, providing some understanding about the levels of theory necessary for reproducing experimental results. While a majority of our previous dataset is focused on halide perovskites, it has now been extended to chalcogenide perovskites and a variety of double perovskites.
@@
Given that our dataset is well-representative of the perovskite chemical space, it is suitable for training surrogate machine learning (ML) models [ml-review, ml-review-2] which can then be deployed for accelerated prediction and screening across millions of possible compositions. Using the DFT data, we trained regularized greedy forest models to predict the band gap and decomposition energy (which shows whether a compound would readily decompose to alternative binary phases) of any ABX3 halide perovskite. This is different from energy above hull that takes into account all possible combinations of phases (theoretical and experimental). We choose decomposition energy as a possible stability metric because the primary decomposition pathway for inorganic and hybrid perovskites is to their binary phases [decompose], and most perovskite films are synthesized using these known binary precursors [binary]. These properties are central in determining the stability and photovoltaic suitability of the materials. Since we have data from different levels of theory, we trained these models within a multi-fidelity framework [ml-amk-2] such that the source of data (or DFT functional) is part of the input. We used both compositional information and well-known elemental or molecular properties as part of the descriptors and achieved prediction accuracy of $\sim$ 98% for the properties of interest. We used K-fold cross-validation to evaluate model overfitting. Figure S1 plots the training and testing RMSE values for decomposition energy and band gap across different folds, showing a consistency in prediction and generally no overfitting.
@@
The best regression models were used to make on-demand predictions for thousands of enumerated ABX3 compounds across the four prototype phases, leading to the discovery of (a) novel stable high-efficiency single-junction solar absorber candidates [ml-amk-2], and (b) novel candidates with high predicted solar-to-hydrogen (STH) efficiencies for photocatalytic water splitting [ml-amk-1]. The latter work utilized empirically predicted electronic band edges to place them relative to H2O redox potential levels. Importantly, the best materials found via DFT+ML screening were also divided into different categories, such as hybrid vs purely inorganic, and Pb-free vs Pb-containing compounds. We recently extended this work to structure-based predictions using crystal graph neural network models [gnn-1, gnn-2, gnn-3]. By accumulating a massive perovskite dataset of $\sim$ 30,000 structures and their DFT-computed energies, atomic forces, and stresses from across all our recent publications, we “fine-tuned” the M3GNET model [m3gnet] to develop a comprehensive interatomic potential that is an effective DFT surrogate for our chosen perovskite chemical space [perovsiap]. Our rigorously optimized “Perovs-IAP” model makes accurate energy and force predictions for bulk, surface slab, and defect-containing perovskite structures, and is capable of rapid geometry optimization of any such structure.
@@
Thus, we now have a rich perovskite dataset and versatile descriptor-based and structure-based predictive models. However, DFT predictions of the decomposition energy or energy above hull are not entirely indicative of the likelihood of synthesis of any novel compound. In the past, the Goldschmidt tolerance factor [gs] and octahedral factor [octahedra] have been used to determine the general suitability of constituent ions to form a structurally robust perovskite based on their ionic radii, but satisfying these factors does not necessarily translate to a compound being synthesizable, especially when it comes to complex perovskite alloys with a distribution of anions and cations. DFT energies are often used for evaluating formability, but are clearly insufficient because calculations are performed at 0 K and do not account for entropic stabilization and disorder unless specifically included. What comes to the rescue here is the existence of many perovskite compounds that were actually synthesized and reported in the literature, and may inform the synthesis likelihood of other similar materials identified via computations.
@@
Classification techniques could be used for synthesis predictions provided the availability of sufficient positive and negative labels, but since studies only report positive examples, typical classifiers are not applicable. As material properties are emergent, if we can map their chemical descriptors to a synthesis label, we could, in theory, predict the order of synthesizability of different compositions. However, this requires a semi-supervised learning approach since it may be nearly impossible to classify a material as definitely non-synthesizable. Publications seldom report compositions that were attempted but failed to materialize in the lab. This means one could only apply positive labels to some of the materials and unknown labels to the remaining materials, which have not been synthesized to date but could be in the future. One technique that has recently been used by researchers for synthesis predictions is “Positive-Unlabeled (PU) learning” [pu-learning, syncotrain, synthesizability-review], where the classifier learns from the positive examples and assigns a probability to the unlabeled points based on similarity of the descriptors. PU-learning has been theoretically validated [pu-theoretical-1, pu-theoretical-2] for its capacity to extract information from an unlabeled set, even under the constraint of accessing only positive-labeled examples and no negative labels.
@@
Frey et al. [mxene] demonstrated the use of PU learning in predicting synthesis scores for MXenes, a novel class of layered materials. They used a decision tree classifier with specific MAX (the bulk phase from which MXenes are derived) and MXene features such as the number of layers, in-plane lattice constant, bond lengths, etc., which were combined with DFT-computed features such as formation energies and Bader charges. The labels for the synthesized compounds were obtained from their work on synthesizing MXenes and MAX phases. The model was able to predict 18 new MXene compositions as synthesizable out of 118 unlabeled samples. Another study by Jang et al. [pucgcnn] used the PU learning approach but combined it with atomic features and the crystal-graph convolutional neural network (CGCNN) framework [cgcnn]. This work used the entire Materials Project (MP) [mp] dataset and provided positive synthesis labels to structures having an Inorganic Crystal Structure Database (ICSD) [icsd] tag. Synthesis probabilities were predicted for materials in the Open Quantum Materials Database (OQMD) [oqmd] and the MP-Virtual database; $\sim$ 12% of the total compounds from OQMD and $\sim$ 20% from MP-virtual datasets were found to be synthesizable.
@@
Another study on a similar dataset was performed by Gu et al. [pu-perovs] who trained a graph convolution network based on MEGNet [megnet] using a combination of perovskite data from AFLOW [aflow], Materials Project, and OQMD [oqmd]. Domain-specific transfer learning was performed to make the model accurate for perovskite structures. A recent study by Zhao et al. [mof-pu] utilized a similar approach for Metal Organic Frameworks (MOFs) and used a CGCNN-based classifier model. Compounds were divided based on pre-defined rules, following which their synthesizability scores were predicted. $\sim$ 12% of the compounds from across the unlabeled set were found to be synthesizable. Yet another recent study from Chung et al. [pu-walsh] predicted the synthesizability of ternary oxides using PU learning, with labels applied specifically for oxides that can be synthesized via solid-state methods.
@@
A major consideration is, of course, the collection of synthesis labels for as many compounds as possible. To do this, we used large language models (LLMs) [Llama32] and extracted chemical formulas of ABX3, A2BX6, or A2BB’X6 halide or chalcogenide compounds reported to be synthesized in various publications. We additionally applied labels to compounds based on their existence in the ICSD, ultimately resulting in 239 compounds with positive synthesis labels. This was used along with a dataset of 832 unlabeled compounds simulated using the GGA-PBE functional to train combined models for predicting both DFT properties and the synthesis probability. Based on prediction and screening across thousands of hypothetical compounds, we discovered potentially hundreds of new stable and synthesizable perovskites with promising optoelectronic properties. Our general workflow is presented in Figure 1. In the following sections, we describe the computational data generation, experimental label extraction, descriptors, different ML models, and results of training and prediction over unlabeled samples. We discuss the best candidates thus identified and present an outlook for the future of perovskite discovery.
@@ 2.1.1 Mannodi Group Dataset
[…]
@@
Figure 2 (b) shows the computed decomposition energy (in eV per ABX3 formula unit) plotted against the band gap (in eV) for a total of 851 compounds, from GGA-PBE calculations. While our prior work also reports data from other GGA and HSE06 functionals, we focus on the PBE dataset for this work because of its relatively larger coverage of the perovskite chemical space of interest. Moreover, the decomposition energy and band gap exhibit a strong correlation between the two functionals, as demonstrated in Figure S4. This suggests that synthesizability predictions derived from either functional should yield consistent trends. Even though the majority of the dataset is cubic and HaP compositions, the models we report later perform well for all phases and ChP, VO, and DP compounds. The band gap values range from close to 0 eV to greater than 6 eV, and decomposition energies range from -1.5 eV (indicating high stability) to 1.5 eV. The formula to calculate the decomposition energy is shown in Eq. 1, where $E_{comp}$ stands for the total energy from DFT for the given composition, and the final term accounts for the mixing entropy in alloys. The entropy term helps account for the stabilization induced by the ionic disorder in the perovskite alloys. Although other entropic contributions are ignored here, using the decomposition energy with mixing entropy information as an input descriptor in synthesizability classification helps encode some entropy information into the model. As an example, the decomposition energy for a compound with B-site mixing will be calculated as shown in Eq. 2. These formulas hold for both ABX3 HaP and ChP compounds. For A2BB’X6 double perovskites, we use binaries based on both B and B’ cations. For A2BX6 vacancy ordered double perovskites, the B site has an oxidation state of 4+, and thus we use BX4 as the binary phase. Most of the larger band gaps arise from double perovskites due to the lack of overlap between BX6 octahedra owing to the alternation of B and B’ cations in the structure. Decomposition energies tend to cluster near the 0 eV mark, and we usually consider a range of -0.5 eV to 0.5 eV as stable, keeping in mind theoretical uncertainties and possible entropic contributions. Other computational details can be found in past works [htdft-1, htdft-2, htdft-3].
@@ 2.1.2 Materials Project Data
We also queried the Materials Project (MP) [mp] database to find perovskite structures distinct from the compounds in our dataset and subsequently combined all the compounds into one dataset. We first filtered all possible compounds with general formulas of ABX3, A2BB’X6, and A2BX6, where the anion X is either a halogen or chalcogen. The MP filtering workflow is pictured in Figure 3. This process led to a total of about 900 compounds. Then, we used the Robocrystallographer tool [robocrystallographer] to filter out the compounds that did not contain specific keywords such as “perovskite”, “perovskite-derived”, and “octahedral”, following which a manual examination was performed to verify that the parsed compound is indeed a perovskite structure. This finally led to 225 compounds, out of which 111 were already present in the Mannodi group dataset. For the 114 unique compounds, we performed additional DFT computations using the exact same GGA-PBE input parameters and convergence criteria used in our dataset, thus obtaining their band gaps and decomposition energies. Subsequently, compounds with zero band gaps were excluded from the training set to ensure model stability, resulting in 58 new compounds which were added to our original dataset to yield an expanded dataset of 909 points. We also used the “ICSD tag” (indicating experimental existence) available in the MP for any of the compounds to add synthesis labels to them. Out of the 225 compounds collected from MP, 33 had a valid ICSD tag, showing that they had been previously successfully synthesized. The computed properties pictured in Figure 2(b) are for the combined PBE dataset of 909 unique compounds.
Figure 3: The workflow used to extract valid perovskite compounds from the MP [mp] database using the Materials Project API
; this data is then combined with our DFT dataset following quick calculations using uniform computational parameters.